Suppose the current exchange rate for the Russian ruble is RUB 34.51. The expected exchange rate in three years is RUB 37.77. What is the difference in the annual inflation rates for the United States and Russia over this period? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

According to relative purchasing power parity equation:

Forward rate = spot rate * (1+( difference in the inflation of foreign country to home country))n

N= period or time

So here,

37.77= 34.51* (1+ x)3

37.77/34.51 = (1+x)3

1.094465 = (1+x)3

1.030545= 1+x

X= 0.03054 or 3.05%

So the difference in the inflation between US and Russia = 3.05%